I am having trouble with this toic can someone please help me.
Solve logarithmic equation
log x + log(x+9) = 1
Thank you inadvance for your help.

Answer by scott8148(6628)

using rules for logarithms, x(x+9)=10 ___ x^2+9x-10=0

factoring gives (x+10)(x-1)=0

REMEMBER : logarithms are NOT defined for negative numbers